If perimeter of rectangle is 40.And the width is 8, what is the lenth?
Answer by mathrebel(25)

Simple: the formula for perimeter of a rectangle is 2*(a+b), where a and b are the sides:
side b=8
side a=?
40=2*(a+8) divide by two both sides and get:
20=a+8 subtract 8 from both sides and:
a=12 whoala!!!
now plug it in back to check the answer:
is 40=2*(12+8)??
40=2*20
40=40 you betcha!!!!
good luck